# Johan Håstad

### v2

@@ Abstract
We study Frege proofs for the functional and onto graph Pigeon Hole Principle defined on the $n\times n$ grid where $n$ is odd. We are interested in the case where each formula in the proof is a depth $d$ formula in the basis given by $\land$, $\lor$, and $\neg$. We prove that in this situation the proof needs to be of size exponential in $n^{\Omega(1/d)}$. If we restrict the size of each line in the proof to be of size $M$ then the number of lines needed is exponential in $n/(\log M)^{O(d)}$. The main technical component of the proofs is to design a new family of random restrictions and to prove the appropriate switching lemmas.
@@ 1 Introduction
[…]
A fundamental and popular case is resolution corresponding to $d=1$, where each formula is a disjunction of literals. It is far from easy to analyze resolution but this proof system has been studied for a long time and many questions are now resolved. We do not want to discuss the history of resolution but as it is very relevant for the current paper let us mention that an early milestone was obtained by Haken [7] in 1985 when he proved that the Pigeon Hole Principle (PHP) requires exponential size resolution proofs. The PHP states that $n+1$ pigeons can fly to $n$ holes such that no two pigeons fly to the same hole. There are many ways to code this statement using Boolean variables and we use the basic variant which has $(n+1)n$ Boolean variables $x_{ij}$. Such a variable is true iff pigeon $i$ flies to the hole $j$. The axioms say that for each $i$ there is a value of $j$ such that $x_{ij}$ is true and for each $j$ there is at most one $i$ such that $x_{ij}$ is true. This is clearly a contradiction but to prove this counting is useful and resolution is not very efficient when it comes to counting.
Previous work. The focus of this paper is the more powerful proof system obtained for larger values of $d$ and here a pioneering result was obtained by Ajtai [1] proving superpolynomial lower bounds for the size of any proof for PHP for any fixed constant $d$. The lower bounds of Ajtai were not explicit and Bellantoni et al [3] gave the first such bounds, namely that depth $\Omega(\log^{*}n)$ is needed for the size of the proof to be polynomial. This was greatly improved in two independent works by Krajíček, Pudlák, and Woods [13] and Pitassi, Beame, and Impagliazzo [16], respectively. These two papers established lower bounds for the size of any proof of the PHP of the form exponential in $n^{c^{-d}}$ where $c>1$, and gave non-trivial bounds for depths as high as $\Theta(\log{\log n})$.
Related questions were studied in circuit complexity where the central question is to study the size of a circuit needed to compute a particular function. Here a sequence of results [6],[20], [23], [8] established size lower bounds of the form exponential in $n^{\Omega(1/d)}$ and obtained strong lower bounds for $d$ as large as $\Theta(\log n/\log{\log n})$. The results were obtained for the parity function and here it is easy to show that this function can be computed by circuits of matching size. To see that PHP allows proof of size exponential in $n^{O(1/d)}$ is more difficult but was established in 2001 by Atserias et al [2].
@@
The first result that broke this barrier was obtained by Pitassi et al [17] who obtained super-polynomial lower bounds for depths up to any $o(\sqrt{\log n})$. The tautology considered was first studied by Tseitin [21] and considers a set of linear equations modulo two defined by a graph. The underlying graph for [17] is an expander. These results were later extended to depth $\Theta(\log n/\log{\log n})$ by Håstad [10] and in this case the underlying graph is the two-dimensional square grid. The bounds obtained were further improved by Håstad and Risse [11].
@@
All results mentioned so far only discuss total size. For resolution, each formula derived is a clause and hence of size at most $n$ but for other proof systems it is interesting to study the number of lines in the proof and the sizes of lines separately. Pitassi, Ramakrishnan and Tan [15] had the great insight that a technical strengthening of the used methods yields much stronger bounds for this measure than implied by the size bounds. They combined some of the techniques of [10] with some methods from [17] to establish that if each line is of size at most $M$ then the number of lines in a proof that establishes the Tseitin principle over the square grid needs to be exponential in $n2^{-d\sqrt{\log M}}$. By using some additional ideas Håstad and Risse [11] fully extended the techniques of [10] to this setting improving the bounds to exponential in $n/(\log M)^{O(d)}$.
@@
Our results. Despite this progress, the thirty year old question whether the PHP allows polynomial size proof of depth $O(\log\log n)$ remained open. The purpose of this paper is to prove that it does not and that lower bounds similar to those for the Tseitin tautology also apply to the PHP. To build on previous techniques we study what is known as the graph PHP where the underlying graph is an odd size two-dimensional grid.
@@
As the side length of the grid is odd, if one colors it as a chess board, the corners are of the same color and let us assume this is white. In the graph PHP on the grid, there is a pigeon on each white square and it should fly to one of the adjacent black squares that define the holes. This graph PHP is the result of the general PHP where most variables are forced to take the value 0. Each pigeon is only given at most four alternatives. Clearly any proof for the general PHP can be modified to give a proof of the graph PHP by replacing some variables by the constant false. To limit ourselves further we prove our lower bounds for what is known as the functional and onto PHP. We add axioms saying that each pigeon can only fly to one hole and each hole receives exactly one pigeon.
@@
Phrased slightly differently, the functional and onto PHP on the grid says that there is a perfect matching of the odd size grid and we heavily use local matchings. We can compare this to the Tseitin tautology on the grid studied by [10, 15, 11] that states that it is possible to assign Boolean values to the edges of the grid such that there is an odd number of true variables next to any node. As a perfect matching would immediately yield such an assignment, the PHP is a stronger statement and possibly easier to refute. In particular, any lower bound for functional and onto PHP gives the same lower bound for the Tseitin contradiction. We do not, however, think of the results of the current paper as a new proofs for the results in [11] (with slightly weaker bounds). In fact, as the construction of this paper shares many features with the construction of [11], it is better to think the current proof as an adjustment of the latter proof taking care of complications that we here only work with matchings. Let us turn to discuss the main technical point, namely to prove a “switching lemma”.
@@
To preserve a tautology or a complicated function it is useful to replace several old variables by the same new variable, possibly negated. It is possible to be even more liberal and allow old variables to be replaced by slightly more complicated expressions in the new variable. This technique was first introduced explicitly by Rossman, Servedio, and Tan [19] when studying the depth hierarchy for small-depth circuits but had been used in more primitive form in earlier papers. We use such generalized restrictions in this paper.
@@
Multi-switching. The technical strengthening needed by [15] that we discussed above is to improve the standard switching lemma to what is commonly known as a multi-switching lemma. This concept was first introduced independently by Håstad [9] and Impagliazzo, Mathews, and Paturi [12] to study the correlation of small-depths circuits and simple functions such as parity.
@@
In this setting one considers many formulas $(F^{i})_{i=1}^{m}$ and the goal is to switch them all simultaneously in the following sense. There is a small depth (common) decision tree such that at any leaf of the tree it is possible to represent each $F^{i}$ by a small formula of the other type. It was the insight that multi-switching could be used in the proof complexity setting that made it possible for [15] to derive the strong bounds on the number of lines in a proof when each line is short. The extension to multi-switching turns out to be mostly technical and not require any really new ideas. In view of this let us for the rest of this introduction only discuss standard switching.
@@
Techniques used. The most novel part is to design this new space of restrictions that preserve the functional and onto PHP. It has many similarities with the space introduced in [10] and from a very high level point of view, the proofs follow the same path. At the more detailed level in this paper we work with partial matchings of the grid which is a more rigid object than assignment that only satisfy the Tseitin condition of an odd number of true variables next to any node. This results in considerable changes in the details and as a result we get slightly worse bounds.
@@
Once the space of restrictions is in place, two tasks remain. Namely to prove the switching lemmas and then use these bounds to derive the claimed bounds on proof size. This latter part hardly changes compared to previous papers. The switching lemmas follow the same pattern as in [10] and [11] but we need some new combinatorial lemmas and the fact that we work with matchings calls for some modifications.
@@
Outline of the paper. We start with some preliminaries and recall some facts from previous papers in Section 2. We introduce our new space of random restrictions in Section 3. The basic switching lemma is proved in Section 4 and we use it to establish the lower bound for proof size in Section 5. We give the multi-switch lemma in Section 6 and use it, in Section 7, to derive the lower bounds on the number of lines in a proof. We end with some very brief comments in Section 8.
@@ 2.1 The formula to refute
We study the functional and onto PHP on the odd size $n\times n$ grid. Nodes are given indexed by $(i,j)$ where $1\leq i,j\leq n$ and a node is connected to other nodes where one of the two coordinates is the same and the other differ by 1. As opposed to some previous papers studying the Tseitin tautology, it is here important that we are on the grid and not on the torus as we want a bipartite graph. For any edge, $e=(v,w)$ of adjacent nodes, $v$ and $w$, we have a variable $x_{e}$. The axioms say that for each $v$ if you consider the, at most four, variables adjacent to $v$ exactly one of them is true. This is easy to write as a 4-CNF. The axioms look exactly the same for holes and pigeons.
This formula is a stronger statement compared to the standard PHP formula with $(n^{2}+1)/2$ pigeons and $(n^{2}-1)/2$ holes. Our parameters are slightly different from the standard parameters where $n$ is the number of holes, but we trust the reader to keep this in mind.
We assume that the grid is colored as a chess board and that the corners are white. Thus pigeons correspond to white nodes and holes to black nodes. We prefer to use the colors to distinguish the two as we later use larger objects that inherit the same colors.
@@ 2.2 Frege proofs
[…]
@@ 2.3 Decision trees and partial matchings of the grid
The use of decision trees is central to the current paper let us define some notions. In a standard decision tree, at each node we ask for the value of an input variable. In this paper we instead allow queries of the form.
@@
Definition 2.1
A partial matching $M$ on the $n\times n$ grid of size $t$ is locally consistent if it can be extended to a complete matching of a larger set $S\times T$. We require that each of $S$ and $T$ is the union of disjoint even-sized intervals such that the total size of all intervals is at most $48t$.
Remark. The constant $48$ in the definition is not optimal and with a more careful analysis it can be decreased, but we have opted for shorter proofs rather than optimal constants.
@@
Lemma 2.2
Suppose we a locally consistent partial matching, $M$ of size at most $n/50-9$ in the $n\times n$ grid and that we are given a node $v$ not matched by $M$. It is then possible to find a partner, $w$, of $v$, such that $M$ jointly with $(u,w)$ is a locally consistent matching.
Proof:  If $v$ is already in $S\times T$ we can use the same extension. Suppose $v=(a,b)$ where $a\in S$ and $b\not\in T$. It is easy to find $b^{\prime}$ such that $T\cup\{b,b^{\prime}\}$ is a union of even size intervals. Now we can add matchings of $S\times b$ and $S\times b^{\prime}$ using that $S$ is a union of even size intervals.
@@
We enlarge the sizes of $S$ and $T$ by at most 2 but since the size of the matching increases by one this is not a problem.
Let us next establish natural property that any sub-matching of a locally consistent matching is locally consistent. With the given definition of locally consistent, the proof needs some work.
Lemma 2.3
@@
Proof:  Suppose the matching $M$ is of size $t$ and let $M_{0}$ be a sub-matching of size $t_{0}$. We have sets of $S$ of $T$ and an extension of $M$ to a matching of $S\times T$. This, of course, is also an extension of $M_{0}$ and if both these sets are of size at most $48t_{0}$ we are done, so assume that this is not the case and we need to decrease the sizes of $S$ and $T$.
Figure 1: A figure that cannot be tiled.
Definition 2.4
The following lemma is one reason why the given definition is useful.
Lemma 2.5
Assume we direct the perimeter as described, i.e. the outer perimeter in the clockwise direction and the perimeter around any holes in the counter-clockwise direction. If the figure contains $b$ black squares and $w$ white squares then the total cost of the perimeter is $4(w-b)$.
Proof:  For each square in the figure consider the four edges of the perimeter of this square when it is traversed in the clockwise order. The total cost of all edges is then $4(w-b)$ as each white square contributes 4 and each black square -4. The edges internal to the figure are traversed once in each direction and these two contributions cancel each other. The edges traversed once give the cost of the perimeter and hence the lemma follows.
Suppose we are given a figure, $A$, with equally many black and white squares and we want to determine whether it admits a matching. As we can treat each connected component separately we can assume that $A$ is connected. We use Hall’s marriage theorem to study this question. Let $L$ be a set of black squares and $N(L)$ we the set of white neighbors, both within $A$. It follows from Hall’s marriage theorem that a necessary and sufficient condition for $A$ to be matchable is that for any $L$ we have $|N(L)|\geq L$. For a set $L$ that potentially violate this condition let us look at the figure $A_{0}=L\cup N(L)$. This contains more black squares than white squares and hence the cost of its perimeter must be negative.
This perimeter contains edges from the perimeter of $A$ and some edges internal to $A$. Note that the latter edges all have cost one as there must be a white square to its right. Indeed if there was a black square to its right, $N(L)$ would contain also the square to its left and hence the edge would not belong to the perimeter. This property implies that these edges internal to $A$ must alternate in the sense that every other edges is vertical and every other edge is horizontal. We call such a sub-figure “white bounded”. Using this terminology we have the following lemma which is a combination of the above reasoning and Lemma 2.5.
Lemma 2.6
For instance in Figure 1 one such sub-figure is bounded by the three white square going down the diagonal in the middle. This sub-figure contains 6 black and 5 white squares.
Let us return to our matching $M_{0}$ which can be extended to a matching of $S\times T$ which are of total size $48t$. The aim is to decrease this size. In order to do this we define a new notion. Let $I=[a,b]$ be an interval and $K$ a multiset contained in the interior of $I$. Define the “left function” of $I$ as follows.
  • $f^{I}_{l}(a)=0$.
Similarly we have a right function, defined by $f^{I}_{r}(b)=0$ and $f^{I}_{r}(i)=f^{I}_{r}(i+1)+\delta_{i}$.
Definition 2.7
We have the following lemma.
Lemma 2.8
Proof:  We add the points of $K$ one by one and build the set of intervals at the same time maintaining the property that the current intervals form a well cover of the so far added points. For each point in $K$ we add at most 5 points to $S$. We make the intervals of even length in the end.
We claim that the resulting set well covers $K$ and establish this by induction. Clearly this is true if $K$ is a singleton and we need to study adding a general point $p$ to $K$.
The claim uses the alternative characterization of $f^{I}_{l}(x)$ as being twice the number of to left of $x$ that belongs to $I\setminus K$ minus half the number of points in $I\cap K$ to the left of or equal to $x$. If there was no merger of interval the bound on $f_{l}^{I}$ is obvious. That mergers do not decrease $f_{l}^{I}$ follows from the fact that the contribution to $f_{l}^{I}$ from $I^{i}$ is non-negative due to the non-negativity of $f_{l}^{I^{i}}$. We conclude that the claim holds.
The squares of $M_{0}$ gives a number of connected figures and consider any point on the perimeter of one of these figures. We define $K_{1}$ to be the set of all first coordinates of these points, considered as a multiset. This is clearly a set of size at most $8t_{0}$. We let $S_{0}$ be the set given by Lemma 2.8 that well covers $K_{1}$. Similarly considering the second coordinates we get a set $K_{2}$ and let $T_{0}$ be a well cover of this set. If either $S_{0}$ or $T_{0}$ contains points outside $[n]$ we drop these point(s). By Lemma 2.8 both $S_{0}$ and $T_{0}$ are of size at most $48t_{0}$ and hence Lemma 2.3 follows from Lemma 2.9 below.
Lemma 2.9
$M_{0}$ can be extended to a matching of $S_{0}\times T_{0}$.
Proof:  Define the figure $A$ to be $S_{0}\times T_{0}$ with the squares of $M_{0}$ removed. Suppose for contradiction that $A$ cannot be tiled with dominos of size 2. Then by Lemma 2.6 we get a white-bounded connected sub-figure, $A_{0}$, with a perimeter of negative cost. As we only analyze $A_{0}$, we only need to consider its connected component in $S_{0}\times T_{0}$ and hence we can assume that $S_{0}$ and $T_{0}$ both are given by single intervals. The perimeter of $A_{0}$ contains an outer boundary and the perimeters of its holes (if any). First we observe that cost of the perimeter of any hole is non-negative. Suppose on the contrary that you have a hole $H$ with negative cost perimeter.
To have any chance to have a negative cost the boundary $H$ hole must contain some elements of $M_{0}$ and these give a matching $M_{0}^{\prime}$. Now consider any rectangle, $R$ with even side lengths that strictly contains $H$. We claim that that there is no extension of $M_{0}^{\prime}$ to cover all of $R$. This follows as the figure obtained by removing from $H$ from $R$ has a negative cost boundary. Furthermore, this figure is of the form $L\cup N(L)$ for a suitable $L$ and its existence would contradict that $M_{0}^{\prime}$ can be extended to any rectangle with even side lengths that contains its component in $S\times T$.
If this boundary contains no part of the boundary of $S_{0}\times T_{0}$ then we can apply the above argument and conclude that the cost is positive. On the other hand if it is identical to the boundary of $S_{0}\times T_{0}$ the cost 0.
In the remaining case we can divide this perimeter into segments $s_{i}$ numbered form 00 to $12t-$ 1 where even numbered segments are along the perimeter of $S_{0}\times T_{0}$ and odd segments contain edges that either are in the interior of $A_{0}$ or along the perimeter of $M_{0}$. First note that $c(s_{2j})\geq 0$ for any $j$. This follows as they start and end with a edge of cost 1. Costs of edges along the perimeter alternate between 1 and -1 except at the corners where we get two adjacent edges of the same cost.
Now let us look at the odd numbered components.
Lemma 2.10
For each odd $i$ we have $c(s_{i})\geq 0$.
Proof:  Suppose that $s_{i}$ starts at point $(a,b)$. Then either $a$ is in the interior of $S_{0}$ or $b$ is in the interior of $T_{0}$ and as the two cases are symmetric we may assume the former. Let us for the time being assume that $b$ is the smallest element in $T_{0}$. If $s_{i}$ does not contain any edges from the perimeter of a component of $M_{0}$ each edge on the segment has cost 1 and the lemma is true. Otherwise let $p$ be the last point of $s_{i}$ that is on the perimeter of $M_{0}$. We claim that the cost of the part of $s_{i}$ from $(a,b)$ to $p$ is non-negative. This establishes the lemma as all edges after $p$ on $s_{i}$ are in the interior of $A_{0}$ and hence of cost 1. Let $b^{\prime}$ be the highest value of any point on the segment from $(a,b)$ to $p$. The following claim, of which we postpone the proof, finishes the proof, in the case when $b$ is the smallest element of $T_{0}$.
Claim 2.11
The path from $(a,b)$ to $p$ has cost at least $f_{l}^{T_{0}}(b^{\prime})$.
The case when $b$ is the largest element in $T_{0}$ is completely analogous. We define $b^{\prime}$ to be the lowest value of the second component of the segment and use $f_{r}^{T_{0}}$ in the corresponding claim. This completes the proof of Lemma 2.10 modulo the claim.
Let us establish Claim 2.11. Look at the projection of the given path on the second coordinate. Consider any $b^{\prime\prime}$ with $b<b^{\prime\prime}\leq b^{\prime}$ such that no point of $M_{0}$ has second coordinate $b^{\prime\prime}$. We want to associate cost two with any such value. As no point in $M_{0}$ projects on $b^{\prime\prime}$ any edge with a second coordinate $b^{\prime\prime}$ must be in the interior of $A_{0}$. As previously observed in this part we alternate between vertical and horizontal edges.
Thus there exists one edge with both end points with second coordinate equal to $b^{\prime\prime}$. We associate the cost 1 to $b^{\prime\prime}$ due to this edge. We must also have two edges with different second coordinates and one end point with second coordinate $b^{\prime\prime}$. In most situation there is one with second coordinate $b^{\prime\prime}-1$ and one with second coordinate $b^{\prime\prime}+1$. The only exception is when $b^{\prime\prime}=b^{\prime}$ and this points does not belong to $M_{0}$ in which case both other and points have second coordinate $b^{\prime\prime}-1$. For each of these two edges we associate cost 1/2 with $b^{\prime\prime}$ for a total cost of two associated with $b^{\prime\prime}$. Note also that the total cost caused by any edge in the interior is 1.
The only edges on the path from $(a,b)$ to $p$ that give value -1 are those from the perimeter of $M_{0}$. We associate the cost -1/2 with each of the second coordinates of its end-points. These associated costs follow closely the definition of $f_{l}^{T_{0}}$ and the claim follows by simple inspection. Having established the claim we conclude that the perimeter of the figure $A_{0}$ cannot have negative costs. This contradiction completes the proof of Lemma 2.9.
@@
We are also interesting in matching areas of the plane which are almost a complete squares. The difference is that some squares next to the perimeter are removed. We call such a removed square a “dent”.
Figure 2: A square with 3 white and 3 black dents´.
Lemma 2.12
@@
Suppose have square with even side length and which has the same number of white and black dents. Suppose further that there are at most $M$ dents and no dent is within $M$ of a corner. In this situation there is always a matching of the square. If we have a square with odd side length with one more white node, then the similar statement is true assuming we have one more white dent.
@@
Proof:  We could use the general condition for the existence of a matching given in Lemma 2.6 but as the situation is very structured we find a direct proof to be easier to follow. It also gives better constants (even though, of course, this is not very important). Note first that as the side lenght is even and we have equally many white and black dents, the number of white squares in the figure equals the number of black squares.

### v3

# On small-depth Frege proofs for $\operatorname{PHP}$

@@ Abstract
We study Frege proofs for the functional and onto graph Pigeon Hole Principle defined on the $n\times n$ grid where $n$ is odd. We are interested in the case where each formula in the proof is a depth $d$ formula in the basis given by $\land$, $\lor$, and $\neg$. We prove that in this situation the proof needs to be of size exponential in $n^{\Omega(1/d)}$. If we restrict each line in the proof to be of size $M$ then the number of lines needed is exponential in $n/(\log M)^{O(d)}$. The main technical component of the proofs is to design a new family of random restrictions and to prove the appropriate switching lemmas.
@@ 1 Introduction
A key property of such a proof system is the kind of statements that can be used and in this paper we study Boolean formulas with at most $d$ alternations. Traditionally one has studied formulas over the basis $\land$, $\lor$, and $\neg$, where negations only appear at the inputs. In this formalism one counts the number of alternations between $\land$ and $\lor$. We instead use the basis given by $\lor$ and $\neg$ and count the number of alternations between the two connectives. There is an easy translation between the two as $\land$ can be simulated by $\neg\lor\neg$. We consider the case where $d$ is a constant or a slowly growing function of the size of the input.
A fundamental and popular case is resolution corresponding to $d=1$, where each formula is a disjunction of literals. It is far from easy to analyze resolution but this proof system has been studied for a long time and many questions are now resolved. We do not want to discuss the history of resolution but as it is very relevant for the current paper let us mention that an early milestone was obtained by Haken [7] in 1985 when he proved that the Pigeon Hole Principle ($\operatorname{PHP}$) requires exponential size resolution proofs. The $\operatorname{PHP}$ states that $n+1$ pigeons can fly to $n$ holes such that no two pigeons fly to the same hole. There are many ways to code this statement using Boolean variables and we use the basic variant which has $(n+1)n$ Boolean variables $x_{ij}$. Such a variable is true iff pigeon $i$ flies to the hole $j$. The axioms say that for each $i$ there is a value of $j$ such that $x_{ij}$ is true and for each $j$ there is at most one $i$ such that $x_{ij}$ is true. This is clearly a contradiction but to prove this counting is useful and resolution is not very efficient when it comes to counting.
Previous work.
The focus of this paper is the more powerful proof system obtained for larger values of $d$ and here a pioneering result was obtained by Ajtai [1] proving superpolynomial lower bounds for the size of any proof for $\operatorname{PHP}$ for any fixed constant $d$. The lower bounds of Ajtai were not explicit and Bellantoni et al [3] gave the first such bounds, namely that depth $\Omega(\log^{*}n)$ is needed for the size of the proof to be polynomial. This was greatly improved in two independent works by Krajíček, Pudlák, and Woods [14] and Pitassi, Beame, and Impagliazzo [15], respectively. These two papers established lower bounds for the size of any proof of the $\operatorname{PHP}$ of the form exponential in $n^{c^{-d}}$ where $c>1$, and gave non-trivial bounds for depths as high as $\Theta(\log{\log n})$.
Related questions were studied in circuit complexity where the central question is to study the size of a circuit needed to compute a particular function. Here a sequence of results [6],[20], [24], [8] established size lower bounds of the form exponential in $n^{\Omega(1/d)}$ and obtained strong lower bounds for $d$ as large as $\Theta(\log n/\log{\log n})$. The results were obtained for the parity function and here it is easy to show that this function can be computed by circuits of matching size. To see that $\operatorname{PHP}$ allows proof of size exponential in $n^{O(1/d)}$ is more difficult but was established in 2001 by Atserias et al [2].
@@
The first result that broke this barrier was obtained by Pitassi et al [17] who obtained super-polynomial lower bounds for depths up to any $o(\sqrt{\log n})$. The tautology investigated was first studied by Tseitin [22] and considers a set of linear equations modulo two defined by a graph. The underlying graph for [17] is an expander. These results were later extended to depth $\Theta(\log n/\log{\log n})$ by Håstad [10] and in this case the underlying graph is the two-dimensional square grid. The bounds obtained were further improved by Håstad and Risse [12].
@@
All results mentioned so far only discuss total size. For resolution, each formula derived is a clause and hence of size at most $n$ but for other proof systems it is interesting to study the number of lines in the proof and the sizes of lines separately. Pitassi, Ramakrishnan, and Tan [16] had the great insight that a technical strengthening of the used methods yields much stronger bounds for this measure than implied by the size bounds. They combined some of the techniques of [10] with methods from [17] to establish that if each line is of size at most $M$ then the number of lines in a proof that establishes the Tseitin principle over the square grid needs to be exponential in $n2^{-d\sqrt{\log M}}$. By using some additional ideas Håstad and Risse [12] fully extended the techniques of [10] to this setting improving the bounds to exponential in $n/(\log M)^{O(d)}$.
Our results.
@@
Despite this progress, the over thirty year old question whether the $\operatorname{PHP}$ allows polynomial size proofs of depth $O(\log\log n)$ remained open. The purpose of this paper is to prove that it does not and that lower bounds similar to those for the Tseitin tautology also apply to the $\operatorname{PHP}$. To build on previous techniques we study what is known as the graph $\operatorname{PHP}$ where the underlying graph is an odd size two-dimensional grid.
@@
As the side length of the grid is odd, if one colors it as a chess board, the corners are of the same color and let us assume this is white. In the graph $\operatorname{PHP}$ on the grid, there is a pigeon on each white square and it should fly to one of the adjacent black squares that define the holes. This graph $\operatorname{PHP}$ is the result of the general $\operatorname{PHP}$ where most variables are forced to take the value 0. Each pigeon is only given at most four alternatives. Clearly any proof for the general $\operatorname{PHP}$ can be modified to give a proof of the graph $\operatorname{PHP}$ by replacing some variables by the constant false. To limit ourselves further we prove our lower bounds for what is known as the functional and onto $\operatorname{PHP}$. We add axioms saying that each pigeon can only fly to one hole and each hole receives exactly one pigeon.
@@
Phrased slightly differently, the functional and onto $\operatorname{PHP}$ on the grid says that there is a perfect matching of the odd size grid and we heavily use local matchings. We can compare this to the Tseitin tautology on the grid studied by [10, 16, 12] that states that it is possible to assign Boolean values to the edges of the grid such that there is an odd number of true variables next to any node. As a perfect matching would immediately yield such an assignment, the $\operatorname{PHP}$ is a stronger statement and possibly easier to refute. In particular, any lower bound for functional and onto $\operatorname{PHP}$ gives the same lower bound for the Tseitin contradiction. We do not, however, think of the results of the current paper as a new proof for the results in [12] (with slightly weaker bounds). In fact, as the construction of this paper shares many features with the construction of [12], it is better to think the current proof as an adjustment of the latter proof taking care of complications that we here only work with matchings. Let us turn to discuss the main technical point, namely to prove a “switching lemma”.
@@
To preserve a tautology or a complicated function it is useful to replace several old variables by the same new variable, possibly negated. It is possible to be even more liberal and allow old variables to be replaced by slightly more complicated expressions in the new variables. This technique was first introduced explicitly by Rossman, Servedio, and Tan [19] when studying the depth hierarchy for small-depth circuits but had been used in more primitive form in earlier papers. We use such generalized restrictions in this paper.
Multi-switching.
@@
The technical strengthening needed by [16] that we discussed above is to improve the standard switching lemma to what is commonly known as a multi-switching lemma. This concept was first introduced independently by Håstad [11] and Impagliazzo, Mathews, and Paturi [13] to study the correlation of small-depths circuits and simple functions such as parity.
@@
In this setting one considers many formulas $(F^{i})_{i=1}^{m}$ and the goal is to switch them all simultaneously in the following sense. There is a small depth (common) decision tree such that at any leaf of the tree it is possible to represent each $F^{i}$ by a small formula of the other type. It was the insight that multi-switching can be used in the proof complexity setting that made it possible for [16] to derive the strong bounds on the number of lines in a proof when each line is short. The extension to multi-switching turns out to be mostly technical and not require any really new ideas. In view of this let us for the rest of this introduction only discuss standard switching.
Techniques used.
@@
The most novel part of this paper is to design a new space of restrictions that preserve the functional and onto graph $\operatorname{PHP}$ on the grid. It has many similarities with the space introduced in [10] and from a very high level point of view, the proofs follow the same path. At the more detailed level in this paper we work with partial matchings of the grid which is a more rigid object than assignments that only satisfy the Tseitin condition of an odd number of true variables next to any node. This results in considerable changes in the details and as a result we get slightly worse bounds.
@@
Once the space of restrictions is in place, two tasks remain. Namely, to prove the switching lemmas and then use these bounds to derive the claimed bounds on proof size. This latter part hardly changes compared to previous papers. The switching lemmas follow the same pattern as in [10] and [12] but we need some new combinatorial lemmas and the fact that we work with matchings calls for some modifications.
Outline of the paper.
@@
We start with some preliminaries and recall some facts from previous papers in Section 2. We introduce our new space of random restrictions in Section 3. The basic switching lemma is proved in Section 4 and we use it to establish the lower bound for proof size in Section 5. We give the multi-switch lemma in Section 6 and use it, in Section 7, to derive the lower bounds on the number of lines in a proof. We end with some very brief comments in Section 8. This is the final version of [9] which was presented at the 2023 FOCS conference.
@@ 2.1 The formula to refute
We study the functional and onto $\operatorname{PHP}$ on the odd size $n\times n$ grid and when we want to emphasize the size of the grid we use the term $\operatorname{PHP}_{n}$. Nodes are given indexed by $(i,j)$ where $1\leq i,j\leq n$ and a node is connected to other nodes where one of the two coordinates is the same and the other differ by $1$. As opposed to some previous papers studying the Tseitin tautology, it is here important that we are on the grid and not on the torus as we want a bipartite graph. For any edge, $e=(v,w)$ of adjacent nodes, $v$ and $w$, we have a variable $x_{e}$. The axioms say that for each $v$ if you consider the, at most four, variables adjacent to $v$ exactly one of them is true. This is easy to write as a 4-CNF. The axioms look exactly the same for holes and pigeons.
This formula is a stronger statement compared to the standard $\operatorname{PHP}$ formula with $(n^{2}+1)/2$ pigeons and $(n^{2}-1)/2$ holes. Our parameters are slightly different from the standard parameters where $n$ is the number of holes, but we trust the reader to keep this in mind.
We assume that the grid is colored as a chess board and that the corners are white. Thus, pigeons correspond to white nodes and holes to black nodes. We prefer to use the colors to distinguish the two as we later use larger objects that inherit the same colors.
@@ 2.2 Frege proofs
We consider proofs where each line in the proof is either an axiom or derived from previous lines. The derivation rules are not important and all we need is that they are of constant size and sound. We use the same rules as [17], [10], [16], and [12]. We demand that each formula that appears is of depth at most $d$ and, as several previous papers, we do not allow $\land$ and count the number of alternations of $\neg$ and $\lor$. The $\land$ operator is simulated by $\neg\lor\neg$. The rules are as follows.
@@ 2.3 Decision trees and partial matchings of the grid
As the use of decision trees is central to the current paper let us define some notions. In a standard decision tree, at each node we ask for the value of an input variable. In this paper we instead allow queries of the form.
@@
Definition 2.1.
A partial matching $M$ on the $n\times n$ grid of size $t$ is locally consistent if it can be extended to a complete matching of a larger set $S\times T$. We require that each of $S$ and $T$ is the union of disjoint even-sized intervals such that the total size of all intervals in each of $S$ and $T$ is at most $48t$.
Remark.
The constant $48$ in the definition is not optimal and with a more careful analysis it can be decreased, but we have opted for shorter proofs rather than optimal constants.
@@
Most of the time we think of parts of the grid as sets of points in $\mathbb{R}^{2}$ with integer coordinates connected in the natural way. For pictures it is sometimes convenient to view them as black and white unit size squares connected if they share and edge. We sometimes use this viewpoint.
Lemma 2.2.
Suppose we have a locally consistent partial matching, $M$ of size at most n/509n/50-9 in the $n\times n$ grid and that we are given a node $v$ not matched by $M$. It is then possible to find a partner, $w$, of $v$, such that $M$ jointly with $(u,w)$ is a locally consistent matching.
Proof 2.3.
If $v$ is already in $S\times T$ we can use the same extension. Suppose $v=(a,b)$ where $a\in S$ and $b\not\in T$. It is easy to find $b^{\prime}$ such that $T\cup\{b,b^{\prime}\}$ is a union of even size intervals. Now we can add matchings of $S\times b$ and $S\times b^{\prime}$ using that $S$ is a union of even size intervals.
@@
We enlarge the sizes of each of $S$ and $T$ by at most 2 but since the size of the matching increases by one this is not a problem.
Let us state the natural property that any sub-matching of a locally consistent matching is locally consistent. As our proof is surprisingly complicated we postpone the proof to Appendix A.
Lemma 2.4.
@@
We are also interested in matching areas of the plane which are almost a complete square. The difference is that some squares next to the perimeter are removed. We call such a removed square a “dent”.
Figure 1: A square with 3 white and 3 black dents.
Lemma 2.5.
@@
Suppose we have a square with even side length and which has the same number of white and black dents. Suppose further that there are at most $M$ dents and no dent is within $M$ of a corner. In this situation there is always a matching of the square. If we have a square with odd side length and white corners, then the similar statement is true assuming we have one more white dent.
Remark 2.6.
The condition of any dent being at distance at least $M$ from the perimeter is not exactly sharp. Some condition on the number of dents next to a corner is, however, needed as it is easy to see that removing many white squares next to a corner gives a local surplus of black squares that cannot be matched. As the exact condition is not important we opted for a simple one.
Proof 2.7.
@@
We could use the general condition for the existence of a matching given in Lemma A.5 but as the situation is very structured we find a direct proof to be easier to follow. It also gives better constants (even though, of course, this is not very important).